Solve the system by addition or substitution.
4x - 4y = -16
y = 5x - 8

Answer by checkley77(12844)

4x - 4y = -16
y = 5x - 8
4X-4(5X-8)=-16
4X-20X+32=-16
-16X=-16-32
-16X=-48
X=-48/-16
X=3 ANSWER.
4*3-4Y=-16
12-4Y=-16
-4Y=-16-12
-4Y=-28
Y=-28/-4
Y=7 ANSWER.
PROOF:
7=5*3-8
7=15-8
7=7